OpenStudy (anonymous):

Simplify the given expression. √-10 divided √-8?

OpenStudy (diyadiya):

\[\frac{\sqrt{-10}}{\sqrt{-8}} =\frac{-\sqrt{2}\sqrt{5}}{-2\sqrt{2}}\] \[=\frac{\sqrt{5}}{2}\]

OpenStudy (anonymous):

srry its multiplication

OpenStudy (diyadiya):

\[\sqrt{-10} \times \sqrt{-8} = -4\sqrt{5}\]